OpenStudy (lgg23):

Trigonometric Integrals and Substitutions: Integral of (tan(x))^(4) dx.

OpenStudy (jamesj):

Well sin^2 x + cos^2 x = 1 Hence tan^2 x + 1 = sec^2 x and tan^2 x = (sec^2 x - 1) Thus first I would write tan^4x = tan^2 x . sec^2 x - tan^2 x = tan^2 x . sec^2 x - (sec^2 x - 1)

OpenStudy (jamesj):

Next you know the derivative of tan x is sec^2 x. Hence you can now integrate the first two of the above three terms.

OpenStudy (jamesj):

Make sense?

OpenStudy (lgg23):

Thanks. I really appreciate it.